### v3

# 3DGStream: On-the-Fly Training of 3D Gaussians for Efficient Streaming of Photo-Realistic Free-Viewpoint Videos

Abstract
Constructing photo-realistic Free-Viewpoint Videos (FVVs) of dynamic scenes from multi-view videos remains a challenging endeavor. Despite the remarkable advancements achieved by current neural rendering techniques, these methods generally require complete video sequences for offline training and are not capable of real-time rendering. To address these constraints, we introduce 3DGStream, a method designed for efficient FVV streaming of real-world dynamic scenes. Our method achieves fast on-the-fly per-frame reconstruction within 12 seconds and real-time rendering at 200 FPS. Specifically, we utilize 3D Gaussians (3DGs) to represent the scene. Instead of the naïve approach of directly optimizing 3DGs per-frame, we employ a compact Neural Transformation Cache (NTC) to model the translations and rotations of 3DGs, markedly reducing the training time and storage required for each FVV frame. Furthermore, we propose an adaptive 3DG addition strategy to handle emerging objects in dynamic scenes. Experiments demonstrate that 3DGStream achieves competitive performance in terms of rendering speed, image quality, training time, and model storage when compared with state-of-the-art methods.
(a) I-NGP [37]: Per-frame training
(b) HyperReel [1]: Offline training
(c) StreamRF [27]: Online training
(d) Ours: Online training
Figure 1: Comparison on the flame steak scene of the N3DV dataset [29]. The training time, requisite storage, and PSNR are computed as averages over the whole video. Our method stands out by the ability of fast online training and real-time rendering, standing competitive in both model storage and image quality.
[…]
1 Introduction
Constructing Free-Viewpoint Videos (FVVs) from videos captured by a set of known-poses cameras from multiple views remains a frontier challenge within the domains of computer vision and graphics. The potential value and application prospects of this task in the VR/AR/XR domains have attracted much research. Traditional approaches predominantly fall into two categories: geometry-based methods that explicitly reconstruct dynamic graphics primitives [14, 16], and image-based methods that obtain new views through interpolation [71, 6]. However, these conventional methods struggle to handle real-world scenes characterized by complex geometries and appearance.
In recent years, Neural Radiance Fields (NeRFs) [34] has garnered significant attention due to its potent capabilities in synthesizing novel views as a 3D volumetric representation. A succession of NeRF-like works  [43, 29, 42, 41, 56, 27, 40, 18, 45, 31, 30, 63] further propelled advancements in constructing FVVs on dynamic scenes. Nonetheless, the vast majority of NeRF-like FVV construction methods encountered two primary limitations: (1) they typically necessitate complete video sequences for time-consuming offline training, meaning they can replay dynamic scenes but are unable to stream them, and (2) they generally fail to achieve real-time rendering, thereby hindering practical applications.
Recently, Kerbl et al. [24] have achieved real-time radiance field rendering using 3D Gaussians (3DGs), thus enabling the instant synthesis of novel views in static scenes with just minutes of training. Inspired by this breakthrough, we propose 3DGStream, a method that utilizes 3DGs to construct Free-Viewpoint Videos (FVVs) of dynamic scenes. Due to procedural reasons, we are currently unable to show the text of our method on arXiv. If you are interested in the implementation, please refer to the source code on our project page or check out the paper published on CVPR Open Access.
3DGStream is capable of rendering photo-realistic FVVs at megapixel resolution in real-time, boasting exceptionally rapid per-frame training speeds and limited model storage requirements. As illustrated in Figs. 1 and 2, compared with static reconstruction methods that train from scratch per-frame and dynamic reconstruction methods that necessitate offline training across the complete video sequences, our approach excels in both training speed and rendering speed, maintaining a competitive edge in image quality and model storage. Furthermore, our method outperforms StreamRF [27], a state-of-the-art technique tackling the exactly same task, in all the relevant aspects.
To summarize, our contributions include:
  • We propose 3DGStream, a method for on-the-fly construction of photo-realistic, real-time renderable FVV on video streams, eliminating the necessity for lengthy offline training on the entire video sequences.
  • We utilize NTC for modeling the transformations of 3DGs, in conjunction with an adaptive 3DG addition strategy to tackle emerging objects within dynamic scenes. This combination permits meticulous manipulation of 3DGs, accommodating scene alterations with limited performance overhead.
  • We conduct extensive experiments to demonstrate 3DGStream’s competitive edge in rendering quality, training time, and requisite storage, as well as its superior rendering speed, compared to existing state-of-the-art dynamic scene reconstruction methods.
2 Related Work
2.1 Novel View Synthesis for Static Scenes
Synthesizing novel views from a set of images of static scenes is a time-honored problem in the domains of computer vision and graphics. Traditional methods such as Lumigraph [21, 7] or Light-Field [26, 47, 9, 15] achieve new view synthesis through interpolation. In recent years, Neural Radiance Fields (NeRF) [34] has achieved photorealistic synthesizing results by representing the radiance field using a multi-layer perceptron (MLP). A series of subsequent works enhance NeRF’s performance in various aspects, such as accelerating training speeds [49, 37, 11, 19, 23, 12], achieving real-time rendering [60, 20, 13, 68, 22, 44], and improving synthesis quality on challenging scenes [33, 2, 3, 52, 35, 4] or sparse inputs [69, 10, 38, 65, 59, 62, 50]. Since the vanilla NeRF employs costly volume rendering, necessitating neural network queries for rendering, subsequent approaches faced trade-offs in training time, rendering speed, model storage, image quality, and applicability. To address these challenges, Kerbl et al. [24] propose 3D Gaussian Splatting (3DG-S), which integrates of 3DGs with differentiable point-based rendering. 3DG-S enables real-time high-fidelity view synthesis in large-scale unbounded scenes after brief training periods with modest storage requirements. Inspired by this work, we extend its application to the task of constructing FVVs of dynamic scenes. Taking it a step further, we design a on-the-fly training framework to achieve efficient FVV streaming.
2.2 Free-Viewpoint Videos of Dynamic Scenes
Constructing FVVs from a set of videos of dynamic scenes is a more challenging and applicable task in the domains of computer vision and graphics. Earlier attempts to address this task pivoted around the construction of dynamic primitives [14, 16] or resorting to interpolation [71, 6]. With the success of NeRF-like methods in novel view synthesis for static scenes, a series of works [58, 70, 28, 64, 30, 31, 51, 39, 41, 43, 48, 1, 57, 29, 18, 45, 8, 42, 53, 27, 55, 32] attempt to use NeRF for constructing FVVs in dynamic scenes. These works can typically be categorized into five types: prior-driven, flow-based, warp-based, those using spatio-temporal inputs, and per-frame training.
Prior-driven methods [58, 70, 28, 64, 25] leverage parametric models or incorporate additional priors, such as skeletons, to bolster performance on the reconstruction of specific dynamic objects, e.g., humans. However, their application is limited and not generalizable to broader scenes.
Flow-based methods [30, 31] primarily focus on constructing FVVs from monocular videos. By estimating the correspondence of 3D points in consecutive frames, they achieve impressive results. Nonetheless, the intrinsic ill-posedness of monocular reconstructions in intricate dynamic scenes frequently calls for supplementary priors like depth, optical flow, and motion segmentation masks.
Warp-based methods [51, 39, 41, 43, 48, 1, 57] assumpt that the dynamics of the scene arise from the deformation of static structures. These methods warp the radiance field of each frame onto one or several canonical frames, achieving notable results. However, the strong assumptions they rely on often prevent them from handling topological variations.
Methods that use spatio-temporal inputs [29, 54, 18, 45, 8, 42, 53] enhance radiance fields by adding a temporal dimension, enabling the querying of the radiance field using spatio-temporal coordinates. While these techniques showcase a remarkable ability to synthesize new viewpoints in dynamic scenes, the entangled scene parameters can constrain their adaptability for downstream applications.
Per-frame training methods [27, 55, 32] adapt to changes in the scene online by leveraging per-frame training, a paradigm we have also adopted. To be specific, StreamRF [27] employs Plenoxels [19] for scene representation and achieves rapid on-the-fly training with minimal storage requirements through techniques like narrow band tuning and difference-based compression. ReRF [55] uses DVGO [49] for scene representation and optimize motion grid and residual grid frame by frame to model inter-frame discrepancies, enabling high-quality FVV streaming and rendering. Dynamic3DG [32] optimizes simplified 3DGs and integrates physically-based priors for high-quality novel view synthesis on dynamic scenes.
Among the aforementioned works, only NeRFPlayer [48], ReRF [55], StreamRF [27], and Dynamic3DG [32] are able to stream FVVs. NeRFPlayer achieves FVV streaming through a decomposition module and a feature streaming module, but it is only able to stream pre-trained models. ReRF and Dynamic3DG are limited to processing scenes with few objects and foreground mask, necessitating minute-level per-frame training times. StreamRF stands out by requiring only a few seconds for each frame’s training to construct high-fidelity FVVs on challenging real-world dynamic scenes with compressed model storage. However, it falls short in rendering speed. Contrarily, our approach matches or surpasses StreamRF in training speed, model storage, and image quality, all while achieving real-time rendering at 200 FPS.
2.3 Concurrent Works
Except for Dynamic3DG, several concurrent works have extended 3DG-S to represent dynamic scenes. Deformable3DG [66] employs an MLP to model the deformation of 3DGs, while [61] introduces a hexplane-based encoder to enhance the efficency of deformation query. Meanwhile, [67, 17] lift 3DG to 4DG primitives for dynamic scene representation. However, these approaches are limited to offline reconstruction and lack streamable capabilities, whereas our work aims to achieve efficient streaming of FVVs with an online training paradigm.
3 Background: 3D Gaussian Splatting
3D Gaussian Splatting (3DG-S) [24] employs anisotropic 3D Gaussians as an explicit scene representation. Paired with a fast differentiable rasterizer, 3DGs achieves real-time novel view synthesis with only minutes of training.
3.1 3D Gaussians as Scene Representation
A 3DG is defined by a covariance matrix $\Sigma$ centered at point (i.e., mean) $\mu$:
$G(x;\mu,\Sigma)=e^{-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)}.$ (1)
To ensure positive semi-definiteness during optimization, the covariance matrix $\Sigma$ is decomposed into a rotation matrix $R$ and a scaling matrix $S$:
$\Sigma=RSS^{T}R^{T}.$ (2)
Rotation is conveniently represented by a unit quaternion, while scaling uses a 3D vector. Additionally, each 3DG contains a set of spherical harmonics (SH) coefficients of to represent view-dependent colors, along with an opacity value $\alpha$, which is used in $\alpha$-blending (Eq. 4).
3.2 Splatting for Differentiable Rasterization
For novel view synthesis, 3DG-S [24] project 3DGs to 2D Gaussian (2DG) splats [72]:
$\Sigma^{\prime}=JW\Sigma W^{T}J^{T}.$ (3)
Here, $\Sigma^{\prime}$ is the covariance matrix in camera coordinate. $J$ is the Jacobian of the affine approximation of the projective transformation, and $W$ is the viewing transformation matrix. By skipping the third row and third column of $\Sigma^{\prime}$, we can derive a $2\times 2$ matrix denoted as $\Sigma_{2d}$. Furthermore, projecting the 3DG’s mean, $\mu$, into the image space results in a 2D mean, $\mu_{2d}$. Consequently, this allows us to define the 2DG in the image space as $G_{2d}(x;\mu_{2d},\Sigma_{2d})$.
Using $\Sigma^{\prime}$, the color $C$ of a pixel can be computed by blending the $N$ ordered points overlapping the pixel:
$C=\sum_{i\in N}c_{i}\alpha_{i}^{\prime}\prod_{j=1}^{i-1}(1-\alpha_{j}^{\prime}).$ (4)
Here, $c_{i}$ denotes the view-dependent color of the $i$-th 3DG. $\alpha_{i}^{\prime}$ is determined by multiplying the opacity $\alpha_{i}$ of the $i$-th 3DG $G$ with the evaluation of the corresponding 2DG $G_{2d}$.
Leveraging a highly-optimized rasterization pipeline coupled with custom CUDA kernels, the training and rendering of 3DG-S are remarkably fast. For instance, for megapixel-scale real-world scenes, just a few minutes of optimization allows 3DGs to achieve photo-realistic visual quality and rendering speeds surpassing 100 FPS.
4 Method
Due to procedural reasons, we are currently unable to show the text of our method on arXiv. If you are interested in the implementation, please refer to the source code on our project page or check out the paper published on CVPR Open Access.
[…]
(dB) (MB) (mins) (FPS)
Static Plenoxels [19] 30.77 4106 23 8.3 $\checkmark$
I-NGP [37] 28.62 48.2 1.3 2.9 $\checkmark$
3DG-S [24] 32.08 47.1 8.3 390 $\checkmark$
Offline DyNeRF [29] 29.58 0.1 260 0.02 $\times$
NeRFPlayer [48] 30.69 17.1 1.2 0.05 $\checkmark$
HexPlane [8] 31.70 0.8 2.4 0.21 $\times$
K-Planes [45] 31.63 1.0 0.8 0.15 $\times$
HyperReel [1] 31.10 1.2 1.8 2.00 $\times$
MixVoxels [53] 30.80 1.7 0.27 16.7 $\times$
Online StreamRF [27] 30.68 17.7/31.4 0.25 8.3 $\checkmark$
Ours 31.67 7.6/7.8 0.20 215 $\checkmark$
Table 1: Quantitative comparison on the N3DV dataset. The training time, required storage and PSNR are averaged over the whole 300 frames for each scene. DyNeRF [29] only report metrics on the flame salmon scene. Considering the initial model.
[…]
(dB) (MB) (mins) (FPS)
Plenoxels [19] 27.15 1015 14 10
I-NGP [37] 28.10 48.2 1.1 4.1
3DG-S [24] 31.31 21.1 2.6 571
StreamRF [27] 26.72 5.7/9.0 0.17 10
Ours 30.79 4.0/4.1 0.10 288
Table 2: Quantitative comparison on the Meet Room dataset. Note that the training time, required storage and PSNR are averaged over the whole 300 frames. Considering the initial model.
5 Experiments
5.1 Datasets
We conduct experiments on two real-world dynamic scene datasets: N3DV dataset [29] and Meet Room dataset [27].
N3DV dataset [29] is captured using a multi-view system of 21 cameras, comprises dynamic scenes recorded at a resolution of 2704$\times$2028 and 30 FPS. Following previous works [29, 8, 45, 27, 48, 53], we downsample the videos by a factor of two and follow the training and validation camera split provided by [29].
Meet Room dataset [27] is captured using a 13-camera multi-view system, comprises dynamic scenes recorded at a resolution of 1280$\times$720 and 30 FPS. Following [27], we utilize 13 views for training and reserved 1 for testing.
5.2 Implementation
We implement 3DGStream upon the codes of 3D Gaussian Splatting (3DG-S) [24], and implement the Neural Transformation Cache (NTC) using tiny-cuda-nn [36]. For the training of initial 3DGs, we fine-tune the learning rates on the N3DV dataset based on the default settings of 3DG-S, and apply them to the Meet Room dataset. For all scenes, we train the NTC for 150 iterations in Stage 1. and train the additional 3DGs for 100 iterations in Stage 2. Please refer to Suppl. for more details.
Figure 3: Comparison of different approaches for modeling the transformation of 3DGs. Conducted on the second frame of the flame salmon video, utilizing identical initial 3DGs.
Figure 4: Comparison of different approaches on the flame salmon scene.
[…]
Baseline 28.39 0
Rnd. Spawn 28.39 971.9
w/o Quant. Ctrl. 28.43 8710.8
Full Model 28.42 477.7
Table 3: Ablation study of the Adaptive 3DG Addition strategy on the flame salmon scene. The metrics are averaged over the whole sequence.
(a) Result of Stage 1
(b) Result of Stage 2
(c) Ground Truth
Figure 5: Quantitative results of the ablation study conducted on the flame steak scene and the coffee martini scene.
5.3 Comparisons
Quantitative comparisons. Our quantitative analysis involves benchmarking 3DGStream on the N3DV dataset and Meet Room dataset, comparing it with a range of representative methods. We take Plenoxels [19], I-NGP [37], and 3DG-S [24] as representatives of fast static scene reconstruction methods, training them from scratch for each frame. StreamRF [27], Dynamic3DG [32], and ReRF [56] are designed for online training in dynamic scenes. Owing to the limitations of Dynamic3DG and ReRF, which necessitate foreground masks and are confined to scenes with fewer objects, and their minute-level per-frame training times, we select StreamRF selected as the representative for online training methods due to its adaptability and training feasibility on the N3DV and MeetRoom datasets. To demonstrate 3DGStream’s competitive image quality, we drew comparisons with the quantitative results reported for the N3DV dataset in the respective papers of DyNeRF [29], NeRFPlayer [48], HexPlane [8], K-Planes [45], HyperReel [1], and MixVoxels [53], all of which are methods for reconstructing dynamic scenes through offline training on entire video sequences.
In Tab. 1, we present the averaged rendering speed, training time, required storage, and peak signal-to-noise ratio (PSNR) over all scenes of the N3DV dataset. For each scene, the latter three metrics are computed as averages over the whole 300 frames. Besides, we provide a breakdown of comparisons across all scenes within the N3DV dataset in the Suppl. To demonstrate the generality of our method, we conducted experiments on the MeetRoom dataset, as introduced by StreamRF [27], and performed a quantitative comparison against Plenoxels [19], I-NGP [37], 3DG-S [24], and StreamRF [27]. The results are presented in Tab. 2. As presented in Tabs. 1 and 2, our method demonstrates superiority through fast online training and real-time rendering, concurrently maintaining a competitive edge in terms of model storage and image quality. Furthermore, among the methods capable of streaming FVVs, our model requires the minimal model storage.
Qualitative comparisons. While our approach primarily aims to enhance the efficiency of online FVV construction, as illustrated in Tabs. 1 and 2, it still achieves competitive image quality. In LABEL:fig:visual_Comparisons, we present a qualitative comparison with I-NGP [37], HyperReel [1], and StreamRF [27] across scenes on the N3DV dataset [29] and the Meet Room dataset [27], with a special emphasis on dynamic objects such as faces, hands, and tongs, as well as intricate objects like labels and statues. It is evident that our method faithfully captures the dynamics of the scene without sacrificing the ability to reconstruct intricate objects. Please refer to our project page for more video results.
Step Overhead (ms) FPS
Render w/o NTC 2.56 390
+ Query NTC 0.62
+ Transformation 0.02
+ SH Rotation 1.46
Total 4.66 215
Table 4: Rendering profilling for the flame salmon scene at megapixel resolution. Note that flame salmon is the most time-consuming to render of all scenes in our experiments.
5.4 Evaluations
Neural Transformation Cache. We utilize distinct approaches to model the transformations of 3DGs from the first to the second frame within the flame salmon video of the N3DV dataset to show the effectiveness of NTC. Fig. 3 shows that, without multi-resolution hash encoding (w/o Hash enc.), the MLP faces challenges in modeling transformations effectively. Additionally, without the warm-up (w/o Warm-up), it takes more iterations for convergence. Besides, even when compared with the direct optimization of the previous frame’s 3DGs (Direct Opt.), NTC demonstrate on-par performance. In Fig. 4, We present the results of different approaches applied across the entire flame salmon video, excluding the first frame (i.e., Frame 0). w/o Hash enc. and w/o Warm-up. are not able to converge swiftly, resulting in accumulating errors as the sequence progresses. Direct Opt. yields the best outcomes but at the cost of inflated storage. Utilizing NTC, in contrast, delivers comparable results with substantially lower storage overhead by eliminating the need for saving all the 3DGs.
Adaptive 3DG Addition. Tab. 3 presents the quantitative results of the ablation study conducted on the flame salmon scene, and more results are presented in Suppl. The base model without Stage 2, and a set of randomly spawned 3DGs (Rnd. Spawn) in equivalent quantities to our spawn strategy, both fail to capture emerging objects. The variant without our quantity control strategy (w/o Quant. Ctrl.) manages to model emerging objects but requires a significantly larger number of additional 3DGs. In contrast, our full model proficiently reconstructs emerging objects using a minimal addition of 3DGs. The ablation study illustrated in Fig. 5 qualitatively showcases the effect of the Adaptive 3DG Addition strategy, highlighting its ability to reconstruct the objects not present in the initial frame, such as coffee in a pot, a dog’s tongue, and flames.
Real-time Rendering. Following 3DG-S [24], we employ the SIBR framework [5] to measure the rendering speed. Once all resources required are loaded onto the GPU, the additional overhead of our approach is primarily the time taken to query the NTC and transform the 3DGs. As detailed in Tab. 4, our method benefits from the efficiency of the multi-resolution hash encoding and the fully-fused MLP [36], which facilitate rapid NTC query. Notably, the most time-consuming step is the SH Rotation. However, our experiments indicate that the SH rotation has a minimal impact on the reconstruction quality, which may be attributed to the 3DGs modeling view-dependent colors through alternative mechanisms (e.g., small 3DGs of varying colors surrounding the object) rather than SH coefficients. Nonetheless, we maintain SH rotation for theoretical soundness.
6 Discussion
The quality of 3DG-S [24] on the initial frame is crucial to 3DGStream. Therefore, we inherit the limitations of 3DG-S, such as high dependence on the initial point cloud. As illustrated in Fig. 5, there are obvious artifacts beyond the windows, attributable to COLMAP’s [46] inability to reconstruct distant landscapes. Hence, our method stands to benefit directly from future enhancements to 3DG-S. Moreover, for efficient on-the-fly training, we limit the number of training iterations, which restricts modeling of drastic motion in Stage 1 and complex emerging objects in Stage 2.
7 Conclusion
We propose 3DGStream, an novel method for efficient Free-Viewpoint Video streaming. Based on 3DG-S [24], we utilizes an effective Neural Transformation Cache to capture the motion of objects. In addition, we propose an Adaptive 3DG Addition strategy to accurately model emerging objects in dynamic scenes. The two-stage pipeline of 3DGStream enables the online reconstruction of dynamic scenes in video streams. While ensuring photo-realistic image quality, 3DGStream achieves on-the-fly training ($\sim$10s per-frame) and real-time rendering ($\sim$200FPS) at megapixel resolution with moderate requisite storage.